# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v2

@@ 2 Methodology for the systematic literature review
Prior to commencing the literature search, some exclusion criteria were defined. First, papers from the current year 2025 were excluded because it has not yet ended, hence the review will only cover the literature up to the end of year 2024. Second, only journal articles were considered, whereas both book chapters and conference papers were excluded. Such a selection is made for two reasons. Whereas it is true that new research directions often emerge from conference papers, which can contain outstanding results, most often the results presented in this form are considered preliminary even by the authors themselves, frequently because felt as lacking the stability required for standard journal publication. In addition, the exclusion of conference papers is made for readability reasons, specifically to ensure that a manageable number of articles can be discussed in full within a limited number of pages. Nevertheless, conference papers that provide useful benchmarks and which are frequently referenced as accurate benchmark models in other journal articles, are anyway referenced throughout the text.
An advanced search query to be submitted to the databases was designed to capture the breadth of the topic while maintaining specificity. Its general form is (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting)  [^1] Specific search queries for the different databases: Scopus: TITLE-ABS-KEY((("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting)) AND PUBYEAR < 2025 AND (EXCLUDE(DOCTYPE,"cr") OR EXCLUDE (DOCTYPE,"cp")). IEEE Xplore: (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting). Filters applied: Journals, 2000 - 2024. Web of Science: (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting). Filters applied: Article, 2000 - 2024. ACM: [[All: "graph neural network"] OR [All: gnn]] AND [All: "time series"] AND [[All: classification] OR [All: forecasting]] AND [E-Publication Date: (01/01/2000 TO 12/31/2024)]. . The first two elements of the query aim to restrict the pool of GNN models to the time series domain, and the second part focuses the search on classification and forecasting applications.
@@
    Duplicate removal: duplicate records were identified and removed.
@@
A graphical representation of this selection process is provided in Fig. 1. A total of 879 records were identified, 355 from Scopus, 67 from IEEE Xplore, 194 from Web of Science, and 263 from ACM. After that, 253 duplicates were removed. Then, the titles and abstracts of the remaining 626 records were manually evaluated to select the pertinent papers. Among these, 333 were discarded because they were considered outside the scope of this review, and 12 were discarded because they were not in English. Hence, 281 records were included in the review.
Figure 1: PRISMA flowchart summarizing the identification and selection process, adapted from [PRISMA_statement]. The flowchart is structured vertically into three key phases: identification (collection of records from databases and removal of duplicates), screening (assessment for eligibility and exclusion of out-of-scope studies), and inclusion (inclusion of papers in the review). The boxes on the right indicate the number of studies excluded at each stage.
As for conference papers, a total of 1,765 records were identified across the four included databases. Assuming the same proportion of duplicates and articles to be excluded as those found for journal papers, about 600 conference papers would remain, making the total number of papers excessive. As previously stated, these conference papers have been excluded from this review. Next, Sec. 3 provides an overview of the selected journal publications.
@@ 3 Overview of the publications
This section provides a comprehensive overview of the 281 selected journal papers. After their selection, the papers were grouped according to their domain of application, as better explained in the thematic analysis of Sec. 5. In the following, some overall bibliographic information about the reviewed articles is presented.
Fig. 2 displays the number of publications over time across different groups. A cumulative positive trend can be observed over time, reflecting an increasing interest in spatio-temporal GNN models across various fields, since 2020. The four most common groups that have seen the greatest increase in publications are “Environment”, “Generic”, “Mobility”, and “Predictive monitoring”. For more details, see also the group numerosity reported in Tab. 2.
Figure 2: Number of journal publications over time across the different groups.
Table 2: Pivot table illustrating the number of publications by year and group.
@@
Energy 1 - 6 6 20 33
Environment - 2 4 21 16 43
Finance - - 3 6 7 16
Generic - 1 4 19 20 44
Health - 1 2 5 7 15
Mobility 2 4 15 15 30 66
Other - 2 4 14 9 29
Predictive monitoring - 2 3 11 19 35
Total 3 12 41 97 128 281
The catalog of the publication sources indicates that, in the studies period, 145 different journals have published articles on GNN models for time series forecasting or classification. Tab. 3 lists the journals with more than 5 published papers.
@@
ACM Trans. Knowl. Discov. Data 10
@@
Applied Intelligence 8
@@
Expert Systems with Applications 7
ACM Trans. Intell. Syst. Technol. 7
@@
IEEE Transactions on Intelligent Transportation Systems 6
IEEE Transactions on Industrial Informatics 6
Applied Energy 6
The diversity of the journals indicates a vast range of potential applications for spatio-temporal GNN models in different fields. This suggests that researchers often select journals based on the specific domain of application. This allows them to reach targeted audiences, and contribute within their respective fields. Such diversity also underlines the interdisciplinary nature of GNNs research, clearly spanning multiple domains. As for the publishers, there are around 40 different publishers, and the ones with more than 5 papers are listed in Tab. 4.
Table 4: Publishers with more than 5 published paper.
Journal Papers
Elsevier 82
Institute of Electrical and Electronics Engineers Inc. 42
Association for Computing Machinery 27
Multidisciplinary Digital Publishing Institute (MDPI) 27
IEEE 18
IEEE Computer Society 13
Springer 10
Springer Science and Business Media 10
John Wiley and Sons Inc 5
In order to examine the degree of collaboration among researchers, a co-authorship network was generated by using the VOSviewer software [VOSviewer]. Fig. 3 depicts authors as nodes of the network, and collaborations as connections, revealing some research clusters. Fig. 3 shows two of these clusters enlarged.
@@
In Fig. 3, the color of the nodes is related to the average publication year of an author, so that the evolution of research groups over time can be evaluated. Two key observations can be made. First, the number of clusters suggests that there are many distinct research groups, indicating a high degree of fragmentation within the community. Second, the color of the nodes indicates that some groups were active only at the beginning and have subsequently disappeared, while others have recently emerged, as indicated by their yellow color. Some authors have had various collaborations over time, as also shown in Fig. 3.
Fig. 4 shows a pie chart of how corresponding authors are distributed among different countries. This visualization provides an overview of the global distribution of research activities. More than the 70% of the authors are affiliated with institutions in China.
@@ 4.2 Overview of graph neural networks
The notion of GNN was initially proposed by Gori et al. in 2005 in Ref. [gori_aggiunto] and by Scarselli et al. in 2009 in Ref. [Scarselli2009] as an architecture designed to implement a function that maps a graph $\mathcal{G}$ with $n$ nodes into an $m$-dimensional Euclidean space. Noticeably, this definition was intended in general for graph data, not specifically time-dependent data, which is the focus of this SLR. In Ref. [Scarselli2009] the target node’s embedding was learned in two steps. First, neighbor information is iteratively propagated with fixed weights within the graph until a stable fixed point is reached (whose existence is guaranteed under certain assumptions by Banach’s fixed point theorem). Second, errors on targets are propagated back across the graph. Notice that in this setup, propagation is performed in “internal time”, not in “clock time”. More in general, the calculation of node embeddings is quite complex, and over time many other methods have been developed. These methods are based on different architectures and sophisticated aggregation mechanisms. Based on these aggregating mechanisms, a taxonomy of GNN models with different classes has been defined. In this review, we adopt a taxonomy for the discussed spatio-temporal GNNs similar to the one proposed by Chen et al. in Ref. [survey_Lucheroni]. More details about that will be given in Subsec. 4.3.
@@ 4.3 Spatio-temporal GNNs and their taxonomy
This specific form of spatio-temporal modeling assumes that the feature information of a node depends on both its own historical values and on the historical data from its neighboring nodes [survey_spatiotemporal]. Notice that once the the graph structure is given, information is propagated (in “internal time”) within it until the output is obtained, as better explained in the following. Our discussion will build up from this perspective.
When it comes to the development of a spatio-temporal GNN model, there are two primary approaches: one that handles spatial and temporal substructures in separate blocks, and another that integrates and processes these two substructures at once [survey_spatiotemporal]. These approaches are implemented by means of modules that are purely spatial, purely temporal, or hybrid combinations of the two. These modules can then be organized into a series of blocks, thus creating the sought spatio-temporal GNN model.
In the taxonomy of spatio-temporal GNNs used in this review, the focus is exclusively on the description of the spatial module that propagates the information within the nodes at fixed time $t$ (not explicitly included in the following equations for simplicity), which, in the reviewed literature, is a fundamental aspect that differentiates one GNN model from another. Based on the mechanism at the basis of the information propagation in the graph at a given time step, spatio-temporal GNNs can be further classified as recurrent GNNs, convolutional GNNs, and attentional GNNs, as discussed in the following. These classes are not mutually exclusive, as some models may fit into hybrid categories. In addition, these classes are not fully exhaustive of the variety of available models, as there may be models that do not align with any of the existing categories. However, for the purpose of this review, the taxonomy discussed above will be considered sufficient.
@@
where $k$ is the index related to iteration (hence running on “internal time”), $f$ is a parametric trainable function, $x_{u}$ is the attribute associated to node $u$, $r_{(u,v)}$ is the weight of the edge $(u,v)$ between the nodes $u$ and $v$, $\mathcal{N}(v)$ denotes the neighborhood of node $v$, and the initial hidden states $h^{(0)}$ are usually randomly initialized. As detailed in the seminal Ref. [Scarselli2009], which introduced recurrent GNNs, in order to ensure the convergence of Eq. (2) it is necessary that the iterative function $f$ is a contraction mapping.
Convolutional GNNs. Convolutional GNNs are closely related to recurrent GNNs. They were introduced by Kipf and Welling in 2017 in Ref. [graph_convolution_kipf]. Unlike recurrent GNNs, which apply the same iterative contraction mapping $f$ until an equilibrium is reached, convolutional GNNs can use different parameters at each updating step, by stacking multiple graph convolutional layers to extract the node embeddings [cit_rev3]. Noticeably, as the number of convolutional layers increases, the convolutional perceptual field expands, and a more abstract representation of information is thus obtained [art_461].
@@
where $f_{k}$ represents the convolutional operation performed at layer $k\in\{1,\dots,K\}$ (with $K$ typically kept small to prevent issues like over-smoothing [oversmoothing]), $\mathcal{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix, $I\in\mathbb{R}^{n\times n}$ is the identity matrix (used to add self-loops in the adjacency matrix $\mathcal{A}$), $\hat{\mathcal{D}}\in\mathbb{R}^{n\times n}$ is the diagonal degree matrix associated to $\hat{\mathcal{A}}$, $H^{(k)}\in\mathbb{R}^{n\times d_{k}}$ are the $d_{k}$-dimensional node embeddings $h^{(k)}_{v}$ of the entire graph produced by the $k^{\text{th}}$ layer, $X\in\mathbb{R}^{n\times d_{0}}$ is the collection of all node attributes $x_{v}$ of the entire graph, and $\Theta_{k}\in\mathbb{R}^{d_{k-1}\times d_{k}}$ is a trainable weight matrix for the $k^{\text{th}}$ layer [survey_Lucheroni]. The mechanism of node embedding generation from Eq. (3) is illustrated in Fig. 6(b), where it is put in contrast with the mechanism of a recurrent GNNs, shown in Fig. 6(a).
@@
Attentional GNNs. In attentional GNNs, the aggregation process uses the attention mechanism [attention] to combine node features. The equations to compute the node embedding $h_{u}$ of layer $k+1$ from the embeddings of layer $k$ for each node $u$ are
@@
Eq. (4a) represents a learnable linear transformation of the node embedding $h_{u}^{(k)}$ using the trainable weight matrix $W_{k}$ of layer $k$. Eq. (4b) computes a pairwise masked attention score between nodes $u$ and $v$. This is done by applying the concatenation operation ||||| | to the vectors $z_{u}^{(k)}$ and $z_{v}^{(k)}$, then calculating a dot product of the concatenation with a learnable weight vector $\vec{a}^{\,(k)}$, and finally applying the LeakyReLU nonlinearity. The quantity $\epsilon_{uv}$ indicates the importance of the features of node $v$ to node $u$. The term “masked” refers to the fact that $\epsilon_{uv}$ are computed for neighboring nodes only. Eq. (4c) normalizes the attention score through the softmax function in order to make the $\alpha_{uv}$ comparable across different nodes. Finally, Eq. (4d) computes the embedding $h_{u}^{(k+1)}$ of the next layer by aggregating the information in neighboring nodes, weighted by the attention scores [GAT]. Graph attention networks can in addition use a multi-head attention mechanism to stabilize the learning process of self-attention. In practice, the node updating operation in graph attention networks is a generalization of the traditional averaging or max-pooling of neighboring nodes, which allows each node to compute a weighted average of its neighbors and identify the most relevant ones [Brody2021HowAA].
@@
The spatial module is responsible for propagating information between nodes, thereby enabling the analysis of cross-sectional inter-dependencies between different variables. As previously discussed in the description of the mechanisms underlying the assumed taxonomy, the spatial module aggregates information starting from immediate neighboring nodes and extends it outward, thereby capturing also the influence of distant nodes. This spatial operation cannot directly account for temporal information. In contrast, the temporal module, often based on architectures such as LSTM, GRU, or self-attention, focuses on the evolution of data over time, independently from cross-sectional node interactions [art_384]. When these different modules are stacked together, spatial and temporal information end up to be used simultaneously, although they are processed in separate modules. This occurs because the computation of each node state is influenced by both spatial and temporal information. This combined approach allows the spatio-temporal GNN model to be a powerful architecture able to effectively capture at the same time both spatial relationships and temporal dynamics.

### v4

@@ 2 Methodology for the systematic literature review
Prior to commencing the literature search, some exclusion criteria were defined. First, papers from the current year 2025 were excluded because it has only just begun, hence the review will only cover the literature up to the end of year 2024. Second, only journal articles published in Q1 and Q2 journals, as ranked by the Scimago Journal Rank, and conference papers from A and A venues, according to the ICORE conference rankings, were included. This selection was made to ensure that the analysis focused exclusively on high-quality, peer-reviewed research published in top-tier venues. Additionally, book chapters, workshops, and tutorials were excluded.
An advanced search query to be submitted to the databases was designed to capture the breadth of the topic while maintaining specificity. Its general form is (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting) Specific search queries for the different databases: Scopus: TITLE-ABS-KEY((("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting)) AND PUBYEAR < 2025. IEEE Xplore: (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting). Filters applied: 2000 - 2024. Web of Science: (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting). Filters applied: 2000 - 2024. ACM: [[All: "graph neural network"] OR [All: gnn]] AND [All: "time series"] AND [[All: classification] OR [All: forecasting]] AND [E-Publication Date: (01/01/2000 TO 12/31/2024)]. . The first two elements of the query aim to restrict the pool of GNN models to the time series domain, and the second part focuses the search on classification and forecasting applications.
@@
    Duplicate removal and venue filtering: duplicate records were removed, along with publications from journals outside Q1–Q2 and conferences not ranked A or A.
@@
A graphical representation of this selection process is provided in Fig. 2. A total of 2663 records were identified, 766 from Scopus, 258 from IEEE Xplore, 292 from Web of Science, and 1347 from ACM. After that, 1618 duplicates and publications from non top-tier venues were removed. Then, the titles and abstracts of the remaining 1045 records were manually evaluated to select the pertinent papers. Among these, 667 were discarded because they were considered outside the scope of this review, and 12 were discarded because they were not in English. As a result, 366 records were included in the review, comprising 262 journal papers and 104 conference papers. The complete list of selected paper can be found in Appendix A and in the accompanying GitHub repository. Next, Sec. 3 provides an overview of the selected publications.
Figure 2: PRISMA flowchart summarizing the identification and selection process, adapted from [302]. The flowchart is structured vertically into three key phases: identification (collection of records from databases, and removal of duplicates and non-top-tier venues), screening (assessment for eligibility and exclusion of out-of-scope studies), and inclusion (inclusion of papers in the review). The boxes on the right indicate the number of studies excluded at each stage.
@@ 3 Overview of the publications
This section provides a comprehensive overview of the 366 selected papers. After their selection, the papers were grouped according to their domain of application, as better explained in the thematic analysis of Sec. 5. In the following, some overall bibliographic information about the reviewed articles is presented.
Fig. 3 presents a stacked column chart showing the number of journal and conference publications over time, broken down by research group. The height of each column represents the total number of publications per year, while the colored segments represent contributions from each group. A clear upward trend can be observed over time, especially in journal papers, reflecting an increasing interest in spatio-temporal GNN models across various fields, since 2020. The groups that experienced the most significant increase are “Environment”, “Generic”, “Mobility”, and “Predictive monitoring”, among which “Generic” and “Mobility” stand out in terms of both number of publications and the novelty of their contributions. For more details, see also the group numerosity reported in Tab. 2.
Figure 3: Number of journal and conference papers over time across the different groups.
Table 2: Pivot table illustrating the number of publications by year and group (journal and conference papers combined).
@@
Energy 1 1 6 5 20 33
Environment 0 2 4 22 15 43
Finance 0 0 5 6 3 14
Health 1 3 5 7 7 23
Mobility 11 6 24 27 34 102
Predictive monitoring 0 2 4 12 22 40
Generic 4 7 11 28 27 77
Other 0 4 5 14 11 34
Total 17 25 64 121 139 366
The catalog of the publication sources indicates that, in the studies period, 131 different Q1 and Q2 journals have published articles on GNN models for time series forecasting or classification. Tab. 3 lists the journals with more than 5 published papers.
@@
ACM Transactions on Knowledge Discovery from Data 9
@@
ACM Transactions on Intelligent Systems and Technology 7
Expert Systems with Applications 7
@@
Applied Energy 6
IEEE Transactions on Industrial Informatics 6
IEEE Transactions on Intelligent Transportation Systems 6
@@
The diversity of the journals indicates a vast range of potential applications for spatio-temporal GNN models in different fields. This suggests that researchers often select journals based on the specific domain of application. This allows them to reach targeted audiences, and contribute within their respective fields. Such diversity also underlines the interdisciplinary nature of GNNs research, clearly spanning multiple domains. As for ranked A* and A conference venues, 18 different venues have hosted publications on GNN models for time series forecasting or classification. The conference venues with 5 or more papers are listed in Tab. 4.
Table 4: Conference venues with more than 5 published paper.
Conference venue Papers
ACM Int. Conf. on Information and Knowledge Management (CIKM) 30
ACM Int. Conf. on Knowledge Discovery and Data Mining (KDD) 18
ACM Int. Conf. on Advances in Geographic Information Systems (SIGSPATIAL) 11
Advances in Neural Information Processing Systems (NeurIPS) 7
Int. Conf. on Machine Learning (ICML) 5
Fig. 4 shows a pie chart of how corresponding authors are distributed among different countries. This visualization provides an overview of the global distribution of research activities. Nearly 70% of the authors are affiliated with institutions based in China.
In order to examine the degree of collaboration among researchers, a co-authorship network was generated by using the VOSviewer software [373]. In this network, authors are represented as nodes, and co-authorships as connections, revealing clusters of researchers who frequently collaborate. Fig. 5 shows the complete network, while Fig. 5 provides an enlarged view of two selected clusters.
@@
In the figures, the color of each node corresponds to the author’s average publication year, allowing the temporal evolution of research groups to be visualized. From the complete network in Fig. 5, two main insights emerge. First, the large number of clusters indicates a highly fragmented research community, with many small, relatively independent groups. Second, the color of the nodes indicates that some groups were active primarily in earlier years and have since faded, while others have emerged more recently, as indicated by the yellow color. Additionally, several authors appear in multiple clusters, reflecting collaborations that span different groups over time, as also shown in Fig. 5.
Additional figures and analyses of trends in the collected papers are available on our GitHub repository.
@@ 4.2 Overview of graph neural networks
The notion of GNN was initially proposed by Gori et al. in 2005 in Ref. [107] and by Scarselli et al. in 2009 in Ref. [332] as an architecture designed to implement a function that maps a graph $\mathcal{G}$ with $n$ nodes into an $m$-dimensional Euclidean space. Noticeably, this definition was intended in general for graph data, not specifically time-dependent data, which is the focus of this SLR. In Ref. [332] the target node’s embedding was learned in two steps. First, neighbor information is iteratively propagated with fixed weights within the graph until a stable fixed point is reached (whose existence is guaranteed under certain assumptions by Banach’s fixed point theorem). Second, errors on targets are propagated back across the graph. Notice that in this setup, propagation is performed in “internal time”, not in “clock time”. More in general, the calculation of node embeddings is quite complex, and over time many other methods have been developed. These methods are based on different architectures and sophisticated aggregation mechanisms. Based on these aggregating mechanisms, a taxonomy of GNN models with different classes has been defined. In this review, we adopt a taxonomy for the discussed spatio-temporal GNNs similar to the one proposed by Chen et al. in Ref. [35]. More details about that will be given in Subsec. 4.3.
@@ 4.3 Spatio-temporal GNNs and their taxonomy
This specific form of spatio-temporal modeling assumes that the feature information of a node depends on both its own historical values and on the historical data from its neighboring nodes [467]. Notice that once the the graph structure is given, information is propagated (in “internal time”) within it until the output is obtained, as better explained in the following. Our discussion will build up from this perspective.
When it comes to the development of a spatio-temporal GNN model, there are two primary approaches: one that handles spatial and temporal substructures in separate blocks, and another that integrates and processes these two substructures at once [467]. These approaches are implemented by means of modules that are purely spatial, purely temporal, or hybrid combinations of the two. These modules can then be organized into a series of blocks, thus creating the sought spatio-temporal GNN model.
In the taxonomy of spatio-temporal GNNs used in this review, the focus is exclusively on the description of the spatial module that propagates the information within the nodes at fixed time $t$ (not explicitly included in the following equations for simplicity), which, in the reviewed literature, is a fundamental aspect that differentiates one GNN model from another. Based on the mechanism at the basis of the information propagation in the graph at a given time step, spatio-temporal GNNs can be further classified as recurrent GNNs, convolutional GNNs, and attentional GNNs, as represented in Fig. 7 and discussed in the following. These classes are not mutually exclusive, as some models may fit into hybrid categories. In addition, these classes are not fully exhaustive of the variety of available models, as there may be models that do not align with any of the existing categories. However, for the purpose of this review, the taxonomy discussed above will be considered sufficient.
Figure 7: Taxonomy of the spatial component of spatio-temporal GNNs, categorizing models into recurrent, convolutional, and attentional.
@@
where $k$ is the index related to iteration (hence running on “internal time”), $f$ is a parametric trainable function, $x_{u}$ is the attribute associated to node $u$, $r_{(u,v)}$ is the weight of the edge $(u,v)$ between the nodes $u$ and $v$, $\mathcal{N}(v)$ denotes the neighborhood of node $v$, and the initial hidden states $h^{(0)}$ are usually randomly initialized. As detailed in the seminal Ref. [332], which introduced recurrent GNNs, in order to ensure the convergence of Eq. (2) it is necessary that the iterative function $f$ is a contraction mapping.
Convolutional GNNs. Convolutional GNNs are closely related to recurrent GNNs. They were introduced by Kipf and Welling in 2017 in Ref. [190]. Unlike recurrent GNNs, which apply the same iterative contraction mapping $f$ until an equilibrium is reached, convolutional GNNs can use different parameters at each updating step, by stacking multiple graph convolutional layers to extract the node embeddings [423]. Noticeably, as the number of convolutional layers increases, the convolutional perceptual field expands, and a more abstract representation of information is thus obtained [31].
@@
where $f_{k}$ represents the convolutional operation performed at layer $k\in\{1,\dots,K\}$ (with $K$ typically kept small to prevent issues like over-smoothing [105]), $\mathcal{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix, $I\in\mathbb{R}^{n\times n}$ is the identity matrix (used to add self-loops in the adjacency matrix $\mathcal{A}$), $\hat{\mathcal{D}}\in\mathbb{R}^{n\times n}$ is the diagonal degree matrix associated to $\hat{\mathcal{A}}$, $H^{(k)}\in\mathbb{R}^{n\times d_{k}}$ are the $d_{k}$-dimensional node embeddings $h^{(k)}_{v}$ of the entire graph produced by the $k^{\text{th}}$ layer, $X\in\mathbb{R}^{n\times d_{0}}$ is the collection of all node attributes $x_{v}$ of the entire graph, and $\Theta_{k}\in\mathbb{R}^{d_{k-1}\times d_{k}}$ is a trainable weight matrix for the $k^{\text{th}}$ layer [35]. The mechanism of node embedding generation from Eq. (3) is illustrated in Fig. 8(b), where it is put in contrast with the mechanism of a recurrent GNNs, shown in Fig. 8(a).
@@
Attentional GNNs. In attentional GNNs, the aggregation process uses the attention mechanism [375] to combine node features. The equations to compute the node embedding $h_{u}$ of layer $k+1$ from the embeddings of layer $k$ for each node $u$ are
@@
Eq. (4a) represents a learnable linear transformation of the node embedding $h_{u}^{(k)}$ using the trainable weight matrix $W_{k}$ of layer $k$. Eq. (4b) computes a pairwise masked attention score between nodes $u$ and $v$. This is done by applying the concatenation operation $||$ to the vectors $z_{u}^{(k)}$ and $z_{v}^{(k)}$, then calculating a dot product of the concatenation with a learnable weight vector $\vec{a}^{\,(k)}$, and finally applying the LeakyReLU nonlinearity. The quantity $\epsilon_{uv}$ indicates the importance of the features of node $v$ to node $u$. The term “masked” refers to the fact that $\epsilon_{uv}$ are computed for neighboring nodes only. Eq. (4c) normalizes the attention score through the softmax function in order to make the $\alpha_{uv}$ comparable across different nodes. Finally, Eq. (4d) computes the embedding $h_{u}^{(k+1)}$ of the next layer by aggregating the information in neighboring nodes, weighted by the attention scores [376]. Graph attention networks can in addition use a multi-head attention mechanism to stabilize the learning process of self-attention. In practice, the node updating operation in graph attention networks is a generalization of the traditional averaging or max-pooling of neighboring nodes, which allows each node to compute a weighted average of its neighbors and identify the most relevant ones [20].
@@
The spatial module is responsible for propagating information between nodes, thereby enabling the analysis of cross-sectional inter-dependencies between different variables. As previously discussed in the description of the mechanisms underlying the assumed taxonomy, the spatial module aggregates information starting from immediate neighboring nodes and extends it outward, thereby capturing also the influence of distant nodes. This spatial operation cannot directly account for temporal information. In contrast, the temporal module, often based on architectures such as LSTM, GRU, or self-attention, focuses on the evolution of data over time, independently from cross-sectional node interactions. When these different modules are stacked together, spatial and temporal information end up to be used simultaneously, although they are processed in separate modules. This occurs because the computation of each node state is influenced by both spatial and temporal information. This combined approach allows the spatio-temporal GNN model to be a powerful architecture able to effectively capture at the same time both spatial relationships and temporal dynamics.